How do you solve:
Write an equation of a line which contains the points (4, -2) and (-1, -2).
Thank you

Answer by palanisamy(496)

The equation of the required line is
[x-4]/(4+1) = (y+2)/(-2+2)
(x-4)/5 = (y+2)/0
0 = y+2
0 = y+2
y+2=0